# A Survey on Computational Pathology Foundation Models: Datasets, Adaptation Strategies, and Evaluation Tasks

### v1

@@ 4.1 Uni-Modal CPathFMs
Uni-modal CPathFMs are generally trained on large, domain-specific pathology datasets using self-supervised contrastive learning frameworks to learn robust representations of pathological images without labeled data. Similar to the development of contrastive learning in natural images, CPathFMs were initially proposed within the MoCo and SimCLR frameworks. Following a transition through the DINO, DINOv2 was established as the leading framework, serving as the foundation for numerous subsequent studies.
@@
DINO-based CPathFMs. As a successful application of SSL on ViT, DINO has been adopted as a framework for training CPathFMs. Campanella et al., Campanella et al. (2024b) compared the performance of DINO and masked autoencoders (MAE) He et al. (2022) on different scales of pathology datasets, ultimately demonstrating the superiority of DINO for pre-training CPathFMs. Kang et al., Kang et al. (2023) focused on domain-aligned pre-training and proposed data augmentation and curation strategies specifically for pathological images.
@@
DINOv2-based CPathFMs. Most studies using the DINOv2 framework, such as UNI Chen et al. (2024), mainly focus on larger ViT models and more extensive and diverse pre-training datasets. Among them, RudolfV Dippel et al. (2024) incorporates the domain knowledge of pathologists in dataset construction. Some methods have adopted pathology-adapted training methods within the DINOv2 framework, such as Aben et al., Aben et al. (2024) who developed the Online Patching method during the pre-training process, allowing for the online high-throughput extraction of patches of arbitrary size and resolution. Additionally, Virchow2 Zimmermann et al. (2024) replaces the original entropy estimator in DINOv2 with the kernel density estimator (KDE). Unlike previous methods, GigaPath focuses on learning the representation of whole-slide images. It first uses the tile-level encoder, trained under the DINOv2 framework, to learn the representation of a sequence of tiles obtained by dividing the whole slide, treating them as visual tokens. These tokens are then fed into LongNet Ding et al. (2023), which utilizes its Dilated Attention mechanism to perform sparse attention computation, thereby obtaining the overall slide representation.
@@
Other Uni-Modal CPathFMs. While the majority of uni-modal methods focus on DINO and DINOv2, some methods employ other SSL frameworks. CTransPath Wang et al. (2022) adds a branch to MoCov3 to generate queries that retrieve semantically similar samples from the memory bank as positive samples, thus guiding the network’s training with a semantically relevant contrastive loss. REMEDIS Azizi et al. (2023) transfers a ResNet model, pre-trained on large-scale natural images, to the SimCLR framework for self-supervised training on pathological images. Additionally, Filiot et al., Filiot et al. (2023) directly train a ViT within the iBOT framework using pathology data.
@@ 4.2 Multi-Modal CPathFMs
Multi-modal CPathFMs enhance the model’s understanding of pathological images by aligning paired image-text data under the visual-language multi-modal SSL frameworks, such as CLIP and CoCa. These methods typically train pre-trained uni-modal modules using uni-modal SSL frameworks before performing joint visual-language pre-training, which has been shown to improve the performance of downstream tasks Lu et al. (2023). Moreover, multi-modality provides an extra perspective for CPathFMs, where pathological visual representations aligned with additional language signals in latent space can assist large language models (LLMs) in understanding pathology knowledge, thereby contributing to the construction of generative foundation AI assistants for pathologists Lu et al. (2024b).
@@
CLIP-based CPathFMs. The success of CLIP on natural images has inspired some works to apply it in the pathology domain. PLIP Huang et al. (2023) and PathCLIP Sun et al. (2024) both fine-tune a pre-trained CLIP model using datasets composed of paired tiles and their captions. PathAsst Sun et al. (2024) leverages PathCLIP as the visual component to build a pathological multi-modal large language model (MLLM). It first trains a fully connected layer between the visual component and the LLM using instructions constructed from question-answer pairs to align the image encoder with the LLM. Then, the LLM is fine-tuned with a small number of instructions. CHIEF Wang et al. (2024) uses an image encoder from CTransPath to encode the tile sequence extracted from WSIs to obtain WSI-level features and encodes anatomical site information as textual features using the original CLIP’s text encoder. The two are combined to obtain rich WSI-level multi-modal representations.
CoCa-based CPathFMs. The multi-modal decoder specifically designed for CoCa enhances the cross-modal capability of CPathFM, which has led to its application in up-to-date CPathFMs. CONCH Lu et al. (2024a) and PRISM Shaikovski et al. (2024) both pre-train an image encoder on pathology datasets using the iBOT and DINOv2 frameworks, respectively, and then further conduct joint visual-language pre-training within the CoCa framework. The difference is that PRISM extends the image encoder to the WSI-level using a Perceiver network Jaegle et al. (2021) and employs WSIs along with their corresponding clinical reports for training. Additionally, PathChat Lu et al. (2024b) is a pathology-focused MLLM, which uses UNI as the image encoder and further adopts a method similar to CONCH for vision-language pre-training. The trained image encoder is then connected to Llama 2 Touvron et al. (2023) and fine-tuned using various instructions.
@@
Building upon prior work, TITAN Ding et al. (2024) develops a multi-modal whole-slide foundation model primarily designed for training a slide encoder. Its pre-training process is divided into three stages. First, a slide encoder is trained using region crops under the iBOT framework with positional encoding incorporated. Subsequently, under the CoCa framework, the slide encoder is trained at both the tile-level and WSI-level, facilitating the gradual development of the model’s ability to comprehend and generate meaningful vision-language representations for WSIs.
@@ 5 Evaluation Tasks
CPathFMs do not target a specific task during the pre-training phase. Instead, a wide range of evaluation tasks are employed after pre-training to assess the model’s ability to extract features from pathology data. These tasks are diverse, and the evaluation tasks for each CPathFM are not standardized, making it challenging to establish a unified benchmark for CPathFMs. Therefore, we provide a summary of the evaluation tasks along with the CPathFMs performing them, as illustrated in Figure 3. We first categorized the evaluation tasks into six major types based on their application objectives, followed by a further subdivision according to their specific objectives (e.g., focusing on tile-level or WSI-level). On this basis, we also considered variations in task settings (e.g., supervised or zero-shot learning). Finally, we summarized which CPathFMs were used to evaluate each type of task.
@@
A series of complex pathology tasks, such as cancer subtyping, biomarker detection, and mutation prediction, are essentially classification problems. As the most commonly used evaluation task, classification has been widely studied in both tile-level and WSI-level under supervised, few-shot, and zero-shot settings. Unlike fully supervised learning, few-shot and zero-shot learning aim to perform on evaluation tasks when the pre-trained model has seen only a small portion or none of the training samples. Classification tasks at the WSI-level are typically weakly supervised, meaning that such tasks only have global annotations (WSI-level) without details of internal regions. If a CPathFM ultimately extracts tile-level features, a WSI-level representation through an aggregator network needs to be obtained when applying this task. In addition to the three common settings mentioned above, some CPathFMs have also evaluated the model’s out-of-distribution generalization ability under settings where there is a distribution shift between the training and testing datasets. Pathology images from different institutions, different staining methods, and different modalities can all contribute to distribution shifts. Additionally, CONCH classified rare diseases to validate its performance on imbalanced data.
Other common task types include retrieval, generation, and segmentation, which also involve both tile-level and WSI-level tasks. Among these, cross-modal retrieval and generation tasks challenge the cross-modal capabilities of CPathFMs. Survival prediction aims to estimate disease-specific survival based on WSIs. Unlike the previous tasks, MLLMs are evaluated through various visual question answering (VQA) tasks, such as close-ended, open-ended, and multiple-choice questions.
Regardless of the evaluation task, evaluation strategies can be broadly categorized into two types: fine-tuning and non-fine-tuning, based on whether the parameters of the pre-trained CPathFMs are updated during the evaluation process. In fine-tuning strategies, the parameters of CPathFMs are updated to adapt to different evaluation tasks. Among non-fine-tuning strategies, the most commonly used approach is linear probing (a.k.a. linear evaluation) in classification scenarios. This method appends a linear classifier to the pre-trained CPathFMs while keeping their parameters frozen during training. It requires less computational resources compared to fine-tuning and is therefore the most common among the surveyed CPathFMs.
In addition to quantitatively analyzable evaluation tasks, some CPathFMs have undergone qualitative analysis. Virchow, RudolfV, and PLIP perform dimensionality reduction and clustering on the representations and observe the results. Lu et al., Lu et al. (2024b) explore the potential of PathChat as an AI copilot specialized for pathology through a follow-up from users in the form of interactive, multi-turn conversations.
@@ 6 Future Directions
Trustworthy CPathFMs ensures fairness, explainability, security, and transparency. Fairness is particularly critical, as predicted outcomes should be independent of sensitive attributes such as race and gender to prevent potential biases in clinical applications. Enhancing the explainability of CPathFMs is also essential to gaining the trust of pathologists and clinicians, as deep learning models often operate as black boxes. Furthermore, addressing security vulnerabilities in CPathFMs, such as adversarial attacks, is necessary to prevent manipulation of model predictions. Finally, transparency in model development, dataset curation, and evaluation procedures is crucial for reproducibility and regulatory approval, ensuring CPathFMs can be safely deployed in clinical workflows.
@@
Development of WSI-Level MLLMs for Pathology VQA. A WSI-level MLLM would allow context-aware analysis of entire whole-slide images while integrating clinical reports, pathology captions, and other textual information. This could significantly improve AI-assisted diagnostics, enabling models to generate pathology reports, answer clinician queries, and assist in complex diagnostic decision-making.
@@ 7 Conclusion
Computational pathology foundation models have emerged as a powerful approach for analyzing histopathological data, playing a crucial role in developing robust and clinically applicable AI-driven pathology solutions. This survey provides a comprehensive review of existing computational pathology foundation models, examining challenges in pre-training datasets, adaptation strategies, and evaluation tasks, while offering a comparative analysis of their strengths and limitations. Finally, we identify key research gaps and propose potential directions for future advancements.
Acknowledgments
G.W. is supported in part by the National Cancer Institute of the National Institutes of Health under Award Number K99CA286966. Y.R.S. is supported in part by the National Institute of Arthritis and Musculoskeletal and Skin Diseases of the National Institutes of Health under Award Number K23AR080791, the Department of Defense under Award Number W81XWH2110819, and the Melanoma Research Alliance Young Investigator Award.

### v2

@@ 4.1 Uni-Modal CPathFMs
Uni-modal CPathFMs are generally trained on large, domain-specific pathology datasets using SSL frameworks to learn robust representations of pathological images without labeled data. Although there are some MIM-based methods, self-supervised contrastive learning methods play a dominant role. Similar to the development of contrastive learning in natural images, CPathFMs were initially proposed within the MoCo Chen et al. (2021) and SimCLR Chen et al. (2020) frameworks. Following a transition through the DINO, DINOv2 was established as the leading framework, serving as the foundation for numerous subsequent studies.
@@
DINO-based CPathFMs. As a successful application of SSL on ViT, DINO has been adopted as a framework for training CPathFMs. Campanella et al., Campanella et al. (2024b) compared the performance of DINO and MAE on different scales of pathology datasets, ultimately demonstrating the superiority of DINO for pre-training CPathFMs. Kang et al., Kang et al. (2023) focused on domain-aligned pre-training and proposed data augmentation and curation strategies specifically for pathological images.
@@
DINOv2-based CPathFMs. Most studies using the DINOv2 framework, such as UNI Chen et al. (2024), mainly focus on larger ViT models and more extensive and diverse pre-training datasets. Among them, RudolfV Dippel et al. (2024) incorporates the domain knowledge of pathologists in dataset construction. Some methods have developed pathology-adapted training methods within the DINOv2 framework, such as Kaiko Aben et al. (2024) which develops the Online Patching method during the pre-training process, allowing for the online high-throughput extraction of patches of arbitrary size and resolution. Virchow2 Zimmermann et al. (2024) replaces the original entropy estimator in DINOv2 with the kernel density estimator (KDE) and PLUTO Juyal et al. (2024) enhances the DINOv2 loss by incorporating a MAE objective and a Fourier-based loss term to regulate the retention of low- and high-frequency components. Additionally, GPFM Ma et al. (2024) selects multiple existing CPathFMs as expert models and builds a unified knowledge distillation framework by performing Expert Knowledge Distillation on top of DINOv2. Unlike previous methods, GigaPath Xu et al. (2024) focuses on learning the representation of whole-slide images. It first uses the tile-level encoder, trained under the DINOv2 framework, to learn the representation of a sequence of tiles obtained by dividing the whole slide, treating them as visual tokens. These tokens are then fed into LongNet Ding et al. (2023), which utilizes its Dilated Attention mechanism to perform sparse attention computation, thereby obtaining the overall slide representation.
@@
Other Uni-Modal CPathFMs. While the majority of uni-modal methods focus on DINO and DINOv2, some methods employ other SSL frameworks. CTransPath Wang et al. (2022b) adds a branch to MoCov3 to generate queries that retrieve semantically similar samples from the memory bank as positive samples, thus guiding the network’s training with a semantically relevant contrastive loss. REMEDIS Azizi et al. (2023) transfers a ResNet model, pre-trained on large-scale natural images, to the SimCLR framework for self-supervised training on pathological images. Additionally, Phikon Filiot et al. (2023) and BEPH Yang et al. (2024) directly train a ViT model within the MIM-based SSL framework iBOT and BEiTv2, respectively.
@@ 4.2 Multi-Modal CPathFMs
Multi-modal CPathFMs enhance the model’s understanding of pathological images by aligning paired image-text data under the visual-language multi-modal SSL frameworks, such as CLIP and CoCa. These methods typically train pre-trained uni-modal modules using uni-modal SSL frameworks before performing joint visual-language pre-training, which has been shown to improve the performance of downstream tasks Lu et al. (2024).
@@
CLIP-based CPathFMs. The success of CLIP on natural images has inspired research aimed at applying it in the pathology domain, where it enhances model interpretability for pathology AI by leveraging paired histopathological images and textual descriptions (e.g., pathology reports and expert annotations). PLIP Huang et al. (2023), PathCLIP Sun et al. (2024) and QuiltNet Ikezogwo et al. (2024) all fine-tune a pre-trained CLIP model using datasets composed of paired tiles and their captions. CHIEF Wang et al. (2024) uses an image encoder from CTransPath to encode the tile sequence extracted from WSIs to obtain WSI-level features and encodes anatomical site information as textual features using the original CLIP’s text encoder. The two are combined to obtain rich WSI-level multi-modal representations. To integrate domain-specific knowledge, Zhou et al., Zhou et al. (2024b) constructed a pathology knowledge graph (KG) and encode it using a knowledge encoder, which then guides visual-language pretraining. Similarly, KEEP Zhou et al. (2024a) builds a disease KG for encoding and employs knowledge-guided dataset structuring to generate tile-caption pairs for pretraining within the CLIP framework, incorporating positive mining, hardest negative sampling, and false negative elimination strategies.
@@
CoCa-based CPathFMs. The multi-modal decoder specifically designed for CoCa enhances the cross-modal capabilities of CPathFM, contributing to its application in the latest CPathFMs. CONCH Lu et al. (2024) and PRISM Shaikovski et al. (2024) both pre-train an image encoder on pathology datasets using the iBOT and DINOv2 frameworks, respectively, and then further conduct joint visual-language pre-training within the CoCa framework. The difference is that PRISM extends the image encoder to the WSI-level using a Perceiver network Jaegle et al. (2021) and employs WSIs along with their corresponding clinical reports for training. Additionally, MUSK Xiang et al. (2025) first trains the image and text encoders separately on unpaired pathology images and text tokens using masked data modeling within the BEiT-3 framework. Then, it aligns the two modalities under the CoCa framework. Building upon prior work, TITAN Ding et al. (2024) develops a multi-modal whole-slide foundation model primarily designed for training a slide encoder. Its pre-training process is divided into three stages. First, a slide encoder is trained under the iBOT framework with positional encoding incorporated. Subsequently, under the CoCa framework, the slide encoder is trained at both the tile-level and WSI-level, facilitating the gradual development of the model’s ability to comprehend and generate meaningful vision-language representations for WSIs.
In addition to the methods investigated in this work, there are other approaches that focus on multi-modal data beyond text and images, such as chest X-ray images Zhang et al. (2023), genetic sequences Jaume et al. (2024), etc., which fall outside the scope of this survey focused on histopathology staining images. Moreover, some works focused on developing multi-modal large language models (MLLMs) as generative foundation AI assistants for pathologists Sun et al. (2024) are not included, as these works typically align multi-modal CPathFMs mentioned above with existing LLMs. The emphasis of these models is on enhancing VQA capabilities rather than training a general feature extractor.
@@ 5 Evaluation Tasks
CPathFMs do not target a specific task during the pre-training phase. Instead, a wide range of evaluation tasks are employed after pre-training to assess the model’s ability to extract features from pathology data. These tasks are diverse, and the evaluation tasks for each CPathFM are not standardized, making it challenging to establish a unified benchmark for CPathFMs. Therefore, we provide a summary of the evaluation tasks along with the CPathFMs performing them, as illustrated in Figure 3. We first categorized the evaluation tasks into six major perspectives based on their application objectives, followed by a further subdivision according to their specific objectives (e.g., focusing on tile-level or WSI-level). On this basis, we also considered variations in task settings (e.g., supervised or zero-shot learning). Finally, we summarized which CPathFMs were used to evaluate each type of task.
@@
A series of complex pathology tasks, such as cancer subtyping, biomarker detection, and mutation prediction, are essentially classification problems. As the most commonly used evaluation task, classification has been widely studied in both tile-level and WSI-level under supervised, few-shot, and zero-shot settings. Unlike fully supervised learning, few-shot and zero-shot learning aim to perform on evaluation tasks when the pre-trained model has seen only a small portion or none of the training samples. Classification tasks at the WSI-level are typically weakly supervised, meaning that such tasks only have global annotations (WSI-level) without details of internal regions. If a CPathFM only extracts tile-level features, a WSI-level representation through an aggregator network needs to be obtained when applying this task. In addition to the three common settings mentioned above, some CPathFMs have also evaluated the model’s out-of-distribution generalization ability under settings where there is a distribution shift between the training and testing datasets. Pathology images from different institutions, different staining methods, and different modalities can all contribute to distribution shifts. Additionally, CONCH classified rare diseases to validate its performance on imbalanced data.
Other task types include retrieval, generation, segmentation, prediction and VQA, which also involve both tile-level and WSI-level tasks. Among these, cross-modal retrieval, generation and VQA tasks challenge the cross-modal capabilities of CPathFMs. In addition to qualitative analysis, some CPathFMs have undergone qualitative analysis. Virchow, RudolfV, BEPH and PLIP perform dimensionality reduction and clustering on the representations and observe the results.
@@ 6 Future Directions
Trustworthy CPathFMs ensures fairness, explainability, security, and transparency. Fairness is especially crucial, as predicted outcomes should be independent of sensitive attributes, such as race, to avoid potential biases in clinical applications. Enhancing the explainability of CPathFMs is also essential to gaining the trust of pathologists and clinicians, as deep learning models often operate as black boxes. Furthermore, addressing security vulnerabilities in CPathFMs, such as adversarial attacks, is necessary to prevent manipulation of model predictions. Finally, transparency in model development, dataset curation, and evaluation procedures is crucial for reproducibility and regulatory approval, ensuring CPathFMs can be safely deployed in clinical workflows.
@@ 7 Conclusion
This survey provides a review of existing computational pathology foundation models, examining challenges in pre-training datasets, adaptation strategies, and evaluation tasks, while offering a comparative analysis of their strengths and limitations. Finally, we have identified key research gaps and proposed potential directions for future advancements.